# Non-Hermitian Origin of Wannier Localizability and Detachable Topological Boundary States

### v2

@@ Acknowledgements.
We thank Akira Furusaki, Shingo Kobayashi, and Shinsei Ryu for helpful discussion. We appreciate the long-term workshop “Recent Developments and Challenges in Topological Phases” (YITP-T-24-03) held at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. D.N., K. Shiozaki, K. Shimomura, and M.S. are supported by JST CREST Grant No. JPMJCR19T2. K. Shiozaki is supported by JSPS KAKENHI Grant Nos. 22H05118 and 23H01097. K. Shimomura is supported by JST SPRING, Grant No. JPMJSP2110. M.S. is supported by JSPS KAKENHI Grant No. 24K00569. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. 24H00945.
@@
“Non-Hermitian Origin of Wannier Localizability and Detachable Topological Boundary States”
@@ I Detachment of topological boundary states
II From point gap to imaginary line gap
@@
$H_{\theta}\left(k_{x},k_{y}\right)$ $\coloneqq\Gamma U(\theta)\Gamma h\left(k_{x},k_{y}\right)U^{\dagger}(\theta)-\text{i}\sigma_{0}\tau_{0}\quad\left(\Gamma\coloneqq\sigma_{z}\tau_{0}\right),$ (II.1)
@@
$U(\theta)$ $\coloneqq\textrm{exp}\left[\dfrac{\text{i}\pi\theta}{4}\sigma_{y}(\tau_{0}+\tau_{z})\right]\textrm{exp}\left[\dfrac{\text{i}\pi\theta}{4}(\sigma_{0}+\sigma_{z})\tau_{y}\right],$ (II.2)
$h\left(k_{x},k_{y}\right)$ $\coloneqq\dfrac{1}{\mathcal{N}}\left[\sin k_{x}\sigma_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)\sigma_{y}+\text{i}\sin k_{y}\sigma_{0}\right]\dfrac{\tau_{0}+\tau_{z}}{2}+\text{i}\sigma_{0}\dfrac{\tau_{0}-\tau_{z}}{2},$ (II.3)
$\mathcal{N}$ $\coloneqq\sqrt{\sin^{2}k_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)^{2}+\sin^{2}k_{y}}.$ (II.4)
@@
$\Gamma H_{\theta}^{\dagger}\Gamma^{-1}=-H_{\theta}.$ (II.5)
@@
$H_{0}\left(k_{x},k_{y}\right)=\dfrac{1}{\mathcal{N}}\left[\sin k_{x}\sigma_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)\sigma_{y}+\text{i}(\sin k_{y}-\mathcal{N})\sigma_{0}\right]\dfrac{\tau_{0}+\tau_{z}}{2},$ (II.6)
@@
$H_{1}\left(k_{x},k_{y}\right)=\dfrac{\sigma_{0}+\sigma_{z}}{2}\dfrac{\text{i}}{\mathcal{N}}\left[\sin k_{x}\tau_{y}+\left(1-\cos k_{x}-\cos k_{y}\right)\tau_{x}-\sin k_{y}\tau_{z}\right]-\dfrac{\sigma_{0}-\sigma_{z}}{2}\text{i}\tau_{z}-\text{i}\sigma_{0}\tau_{0}.$ (II.7)
Figure S1: Opening of an imaginary line gap while keeping a point gap. The complex spectrum of the point-gapped two-dimensional model $H_{\theta}\left(k_{x},k_{y}\right)$ in class AIII [Eq. (II.1)] is shown for (a) $\theta=0.0$, (b) $\theta=0.3$, (c) $\theta=0.7$, and (d) $\theta=1.0$.
III Bulk-boundary correspondence
Our classification also uncovers the topological nature of the detached boundary states. For a gapped $d$-dimensional Hermitian Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$, the bulk topological invariant is defined [4, 7]. For the semi-infinite boundary conditions with a boundary perpendicular to the $x_{d}$ direction, the boundary Hamiltonian ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ is defined, where ${\bm{k}}_{\parallel}=(k_{1},\dots,k_{d-1})$ is the Bloch momentum for the surface Brillouin zone. Note that the matrix size of ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ is infinite due to the bulk degrees of freedom. Suppose that the boundary Hamiltonian ^$\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ hosts states detached from the bulk spectrum; the orthogonal projection onto the detached states is defined as
$\hat{P}\,({\bm{k}}_{\parallel})=\frac{1}{2\pi\text{i}}\oint_{C}\frac{dz}{z-\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})},$ (III.1)
@@ III Bulk-boundary correspondence
$\hat{\cal C}\hat{P}^{*}\,({\bm{k}}_{\parallel})\,\hat{\cal C}^{-1}$ $=\hat{P}\,(-{\bm{k}}_{\parallel})\quad(\hat{\cal C}\hat{\cal C}^{*}=\pm 1),$ (III.2)
$\hat{\Gamma}\hat{P}\,({\bm{k}}_{\parallel})\,\hat{\Gamma}^{-1}$ $=\hat{P}\,({\bm{k}}_{\parallel})$ (III.3)
with unitary operators $\hat{\cal C}$ and $\hat{\Gamma}$ for particle-hole and chiral symmetries. Accordingly, we define the $(d-1)$-dimensional topological invariants for the detached boundary states protected by $\hat{\cal C}$ and $\hat{\Gamma}$, as specified below.
@@
${\rm Ch}_{n}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]\coloneqq\frac{1}{n!}\left(\frac{\text{i}}{2\pi}\right)^{n}\int_{T^{d-1}}{\rm Tr}\left[\left(\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\,d\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\,d\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right)^{n}\right].$ (III.4)
@@
In the presence of particle-hole symmetry, as $\hat{\cal C}$ behaves as time-reversal symmetry, one can define $\mathbb{Z}_{2}$ invariants for the detached boundary states. Specifically, one can define the class-AII-type $\mathbb{Z}_{2}$ invariant $\nu^{d-1}_{\rm class\ AII}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]$ for class C and the boundary dimensions $d-1=8n+2,8n+3$, and the class-AI-type $\mathbb{Z}_{2}$ invariant $\nu^{d-1}_{\rm class\ AI}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]$ for class D and $d-1=8n+6,8n+7$ [99, 5, 100]. Furthermore, since particle-hole symmetry $\hat{\cal C}$ does not change the chirality (i.e., $\hat{\cal C}\hat{\Gamma}^{*}=\hat{\Gamma}\hat{\cal C}$) for classes CII and BDI, the class-AII-type ${\mathbb{Z}}_{2}$ invariant $\nu^{d-1}_{\rm class\ AII}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]$ for each sector $\hat{\Gamma}=\pm$ is well defined for class CII and $d-1=8n+2,8n+3$. Similarly, the class-AI-type ${\mathbb{Z}}_{2}$ invariant $\nu^{d-1}_{\rm class\ AI}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]$ for each sector $\hat{\Gamma}=\pm$ is well defined for class BDI and $d-1=8n+6,8n+7$.
@@
The imaginary-line-gap to point-gap map described in Sec. IV gives the relationship between the bulk and boundary invariants, summarized as follows. In the presence of chiral symmetry, we have
${\rm Ch}_{n}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]-{\rm Ch}_{n}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.5)
@@
where $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]$ is the $(2n+1)$-dimensional winding number defined by the chiral-symmetric Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$. For classes DIII and CI, since particle-hole symmetry flips chirality as $\hat{\cal C}\hat{\Gamma}^{*}=-\hat{\Gamma}\hat{\cal C}$, we have ${\rm Ch}_{n}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]=-{\rm Ch}_{n}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]$, meaning that if the boundary states are detached, the bulk winding number $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]$ should be even. This accounts for the “$\mathbb{Z}^{\checkmark/\times}$ classification” in Table 1. For the $\mathbb{Z}_{2}$ topological phases in the $d$-dimensional bulk, we obtain
@@
${\rm Ch}_{4n}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+1}_{\rm class\ D}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~{}2\right),$ (III.6)
@@
$\frac{1}{2}{\rm Ch}_{4n}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+1}_{\rm class\ DIII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~{}2\right),$ (III.7)
$\nu^{d-1=8n+2}_{\rm class\ AII}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+2}_{\rm class\ AII}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+3}_{\rm class\ CII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.8)
$\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+4}_{\rm class\ CII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.9)
$\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+4}_{\rm class\ C}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.10)
@@
${\rm Ch}_{4n+2}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+5}_{\rm class\ C}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~{}2\right),$ (III.11)
@@
$\frac{1}{2}{\rm Ch}_{4n+2}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+5}_{\rm class\ CI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~{}2\right),$ (III.12)
$\nu^{d-1=8n+6}_{\rm class\ AI}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+6}_{\rm class\ AI}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+7}_{\rm class\ BDI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.13)
$\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+8}_{\rm class\ BDI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (III.14)
$\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+8}_{\rm class\ D}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right].$ (III.15)
@@
Here, the zeroth Chern number ${\rm Ch}_{0}\,[\hat{P}]$ is the rank of $\hat{P}$. These formulas (III.5)-(III.15) clarify the topology of the detached boundary states, leading to unique boundary responses and disorder effects.
@@
IV Intrinsic non-Hermitian topology

### v3

@@ Acknowledgements.
We thank Akira Furusaki, Zongping Gong, Shu Hamanaka, Shingo Kobayashi, and Shinsei Ryu for helpful discussion. We appreciate the long-term workshops “Recent Developments and Challenges in Topological Phases” (YITP-T-24-03) and “Dynamics Days Asia Pacific 13 (DDAP13) / YKIS2024” held at Yukawa Institute for Theoretical Physics (YITP), Kyoto University. D.N., K. Shiozaki, K. Shimomura, and M.S. are supported by JST CREST Grant No. JPMJCR19T2. D.N. is supported by JSPS KAKENHI Grant No. JP24K22857. K. Shiozaki is supported by JSPS KAKENHI Grants No. JP22H05118 and No. JP23H01097. K. Shimomura is supported by JST SPRING, Grant No. JPMJSP2110 and JSPS KAKENHI Grant No. JP25KJ1632. M.S. is supported by JSPS KAKENHI Grants No. JP24K00569 and No. JP25H01250. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. JP24H00945.
@@
“Non-Hermitian Origin of Detachable Boundary States in Topological Insulators”
@@ I Detachment of topological boundary states
II Point-gap topology and exceptional points
In Fig. S1, we show the complex spectrum of the non-Hermitian model [i.e., Eq. (3) in the main text]
$\tilde{H}_{\rm A}\left(k_{x}\right)=\begin{pmatrix}H_{\rm A}\left(k_{x}\right)&v\\ v^{*}&0\end{pmatrix},\quad H_{\rm A}\left(k_{x}\right)=\sin k_{x}+\text{i}\gamma\left(\cos k_{x}-1\right),$ (II.1)
which describes a non-Hermitian counterpart of the chiral edge state with an additional trivial band. For the sufficiently weak coupling strength $\left|v\right|<\gamma$, a point gap around $E=-\text{i}\gamma$ remains open, accompanied by the nontrivial point-gap topology $W_{1}=1$ (i.e., winding number of the complex spectrum) [32, 41]. Since this point-gap topology is intrinsic to non-Hermitian systems and irreducible to line-gap topology, it necessitates the swapping of the two bands as momentum $k_{x}$ is traversed through the Brillouin zone [see Fig. S1 (b-d)]. Meanwhile, the point gap closes at $\left|v\right|=\gamma$, making the point-gap topology around $E=-\text{i}\gamma$ trivial for $\left|v\right|>\gamma$ [see Fig. S1 (e, f)]. In such a case, this non-Hermitian system can no longer be identified as a chiral edge state.
The swapping of complex energy bands is closely related to exceptional points [Fig. S1 (g)]. In the parameter space $\left(v,k_{x}\right)$, a pair of exceptional points appears at $\left(v,k_{x}\right)=\left(\pm\gamma,\pi\right)$, connected by a branch cut. For $\left|v\right|<\gamma$, traversing the Brillouin zone in momentum space entails crossing this branch cut, resulting in the exchange of the two bands.
Figure S1: Complex spectrum of the non-Hermitian model in Eq. (II.1) for $\gamma=0.8$ and (a) $v=0$, (b) $v=0.05$, (c) $v=0.6$, (d) $v=0.75$, (e) $v=0.8$, and (f) $v=0.85$. The two bands are denoted by blue and magenta. (g) Argument of $E_{+}-E_{-}$ as a function of $v$ and $k_{x}$, where $E_{\pm}\coloneqq(H_{\rm A}\pm\sqrt{H_{\rm A}^{2}+4\,|v|^{2}})/2$ represent the two complex eigenenergies of $\tilde{H}_{\rm A}$. A pair of exceptional points (black crosses) appears at $\left(v,k_{x}\right)=\left(\pm\gamma,\pi\right)$, where the argument is ill defined. These exceptional points are connected through the branch cut, which appears even for $v=0$ owing to the definitions of $E_{\pm}$, involving the square root of the complex function, while the entire spectrum is consistent with $\tilde{H}_{\rm A}$.
III From point gap to imaginary line gap
@@
$H_{\theta}\left(k_{x},k_{y}\right)$ $\coloneqq\Gamma U(\theta)\Gamma h\left(k_{x},k_{y}\right)U^{\dagger}(\theta)-\text{i}\sigma_{0}\tau_{0}\quad\left(\Gamma\coloneqq\sigma_{z}\tau_{0}\right),$ (III.1)
@@
$U(\theta)$ $\coloneqq\textrm{exp}\left[\dfrac{\text{i}\pi\theta}{4}\sigma_{y}(\tau_{0}+\tau_{z})\right]\textrm{exp}\left[\dfrac{\text{i}\pi\theta}{4}(\sigma_{0}+\sigma_{z})\tau_{y}\right],$ (III.2)
$h\left(k_{x},k_{y}\right)$ $\coloneqq\dfrac{1}{\mathcal{N}}\left[\sin k_{x}\sigma_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)\sigma_{y}+\text{i}\sin k_{y}\sigma_{0}\right]\dfrac{\tau_{0}+\tau_{z}}{2}+\text{i}\sigma_{0}\dfrac{\tau_{0}-\tau_{z}}{2},$ (III.3)
$\mathcal{N}$ $\coloneqq\sqrt{\sin^{2}k_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)^{2}+\sin^{2}k_{y}}.$ (III.4)
@@
$\Gamma H_{\theta}^{\dagger}\Gamma^{-1}=-H_{\theta}.$ (III.5)
@@
$H_{0}\left(k_{x},k_{y}\right)=\dfrac{1}{\mathcal{N}}\left[\sin k_{x}\sigma_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)\sigma_{y}+\text{i}(\sin k_{y}-\mathcal{N})\sigma_{0}\right]\dfrac{\tau_{0}+\tau_{z}}{2},$ (III.6)
@@
$H_{1}\left(k_{x},k_{y}\right)=\dfrac{\sigma_{0}+\sigma_{z}}{2}\dfrac{\text{i}}{\mathcal{N}}\left[\sin k_{x}\tau_{y}+\left(1-\cos k_{x}-\cos k_{y}\right)\tau_{x}-\sin k_{y}\tau_{z}\right]-\dfrac{\sigma_{0}-\sigma_{z}}{2}\text{i}\tau_{z}-\text{i}\sigma_{0}\tau_{0}.$ (III.7)
Figure S2: Opening of an imaginary line gap while keeping a point gap. The complex spectrum of the point-gapped two-dimensional model $H_{\theta}\left(k_{x},k_{y}\right)$ in class AIII [Eq. (III.1)] is shown for (a) $\theta=0.0$, (b) $\theta=0.3$, (c) $\theta=0.7$, and (d) $\theta=1.0$.
In Fig. S3, we also show the complex spectrum of the non-Hermitian model [i.e., Eq. (7) in the main text]
$\tilde{H}_{\rm AIII}\left(\bm{k}\right)=\begin{pmatrix}H_{\rm AIII}\left(\bm{k}\right)&v\sigma_{-}\\ v^{*}\sigma_{+}&\varepsilon\sigma_{x}\end{pmatrix},\quad H_{\rm AIII}\left(\bm{k}\right)=\left(\sin k_{x}\right)\sigma_{x}+\left(1-\cos k_{x}-\cos k_{y}\right)\sigma_{y}+\text{i}\gamma\left(\sin k_{y}-1\right),$ (III.8)
which also shows the continuous deformation from point-gap topology to imaginary-line-gap topology.
Figure S3: Complex spectrum of the non-Hermitian model in Eq. (III.8) for $\gamma=1$, $\varepsilon=0.1$, as well as (a) $v=0.0$, (b) $v=0.3$, and (c) $v=0.9$. The Dirac surface state around $E=0$ (orange) can be detached from the bulk bands (blue) in the presence of the perturbation $\varepsilon\sigma_{x}$.
IV Bulk-boundary correspondence
Our classification uncovers the topological nature of the detached boundary states. For a gapped $d$-dimensional Hermitian Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$, the bulk topological invariant is defined [4, 7]. For the semi-infinite boundary conditions with a boundary perpendicular to the $x_{d}$ direction, the boundary Hamiltonian $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ is defined, where ${\bm{k}}_{\parallel}=(k_{1},\dots,k_{d-1})$ is the Bloch momentum for the surface Brillouin zone. Note that the matrix size of $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ is infinite due to the bulk degrees of freedom. Suppose that the boundary Hamiltonian $\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})$ hosts states detached from the bulk spectrum; the orthogonal projection onto the detached states is defined as
$\hat{P}\,({\bm{k}}_{\parallel})=\frac{1}{2\pi\text{i}}\oint_{C}\frac{dz}{z-\hat{\mathcal{H}}_{\rm bdy}\,({\bm{k}}_{\parallel})},$ (IV.1)
@@ IV Bulk-boundary correspondence
$\hat{\cal C}\hat{P}^{*}\,({\bm{k}}_{\parallel})\,\hat{\cal C}^{-1}$ $=\hat{P}\,(-{\bm{k}}_{\parallel})\quad(\hat{\cal C}\hat{\cal C}^{*}=\pm 1),$ (IV.2)
$\hat{\Gamma}\hat{P}\,({\bm{k}}_{\parallel})\,\hat{\Gamma}^{-1}$ $=\hat{P}\,({\bm{k}}_{\parallel})$ (IV.3)
with unitary operators $\hat{\cal C}$ and $\hat{\Gamma}$ for particle-hole and chiral symmetries. Accordingly, we define the (d1)(d-1)-dimensional topological invariants for the detached boundary states protected by $\hat{\cal C}$ and $\hat{\Gamma}$, as specified below.
@@
${\rm Ch}_{n}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]\coloneqq\frac{1}{n!}\left(\frac{\text{i}}{2\pi}\right)^{n}\int_{T^{d-1}}{\rm Tr}\left[\left(\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\,d\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\,d\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right)^{n}\right].$ (IV.4)
@@
In the presence of particle-hole symmetry, as $\hat{\cal C}$ behaves as time-reversal symmetry, one can define $\mathbb{Z}_{2}$ invariants for the detached boundary states. Specifically, one can define the class-AII-type $\mathbb{Z}_{2}$ invariant $\nu^{d-1}_{\rm class\ AII}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]$ for class C and the boundary dimensions d1=8n+2,8n+3d-1=8n+2,8n+3, and the class-AI-type $\mathbb{Z}_{2}$ invariant $\nu^{d-1}_{\rm class\ AI}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]$ for class D and d1=8n+6,8n+7d-1=8n+6,8n+7 [99, 5, 100]. Furthermore, since particle-hole symmetry $\hat{\cal C}$ does not change the chirality (i.e., $\hat{\cal C}\hat{\Gamma}^{*}=\hat{\Gamma}\hat{\cal C}$) for classes CII and BDI, the class-AII-type ${\mathbb{Z}}_{2}$ invariant $\nu^{d-1}_{\rm class\ AII}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]$ for each sector $\hat{\Gamma}$$\pm$ is well defined for class CII and d1=8n+2,8n+3d-1=8n+2,8n+3. Similarly, the class-AI-type ${\mathbb{Z}}_{2}$ invariant $\nu^{d-1}_{\rm class\ AI}\left[\hat{P}^{(\pm)}\,({\bm{k}}_{\parallel})\right]$ for each sector $\hat{\Gamma}$$\pm$ is well defined for class BDI and d1=8n+6,8n+7d-1=8n+6,8n+7.
@@
The imaginary-line-gap to point-gap map described in Sec. V gives the relationship between the bulk and boundary invariants, summarized as follows. In the presence of chiral symmetry, we have
${\rm Ch}_{n}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]-{\rm Ch}_{n}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.5)
@@
where $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]$ is the $(2n+1)$-dimensional winding number defined by the chiral-symmetric Hamiltonian $\mathcal{H}_{\rm bulk}\,({\bm{k}})$. For classes DIII and CI, since particle-hole symmetry flips chirality as $\hat{\cal C}\hat{\Gamma}^{*}=-\hat{\Gamma}\hat{\cal C}$, we have ${\rm Ch}_{n}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]=-{\rm Ch}_{n}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]$, meaning that if the boundary states are detached, the bulk winding number $W_{2n+1}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]$ should be even. This accounts for the $\mathbb{Z}^{\checkmark/\times}$ classification” in Table 1. For the $\mathbb{Z}_{2}$ topological phases in the $d$-dimensional bulk, we obtain
@@
${\rm Ch}_{4n}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+1}_{\rm class\ D}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~2\right),$ (IV.6)
@@
$\frac{1}{2}{\rm Ch}_{4n}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+1}_{\rm class\ DIII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~2\right),$ (IV.7)
$\nu^{d-1=8n+2}_{\rm class\ AII}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+2}_{\rm class\ AII}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+3}_{\rm class\ CII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.8)
$\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+4}_{\rm class\ CII}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.9)
$\nu^{d-1=8n+3}_{\rm class\ AII}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+4}_{\rm class\ C}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.10)
@@
${\rm Ch}_{4n+2}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+5}_{\rm class\ C}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~2\right),$ (IV.11)
@@
$\frac{1}{2}{\rm Ch}_{4n+2}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]\equiv\nu^{d=8n+5}_{\rm class\ CI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right]\quad\left(\mathrm{mod}~2\right),$ (IV.12)
$\nu^{d-1=8n+6}_{\rm class\ AI}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+6}_{\rm class\ AI}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+7}_{\rm class\ BDI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.13)
$\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}^{(+)}\,({\bm{k}}_{\parallel})\right]+\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}^{(-)}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+8}_{\rm class\ BDI}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right],$ (IV.14)
$\nu^{d-1=8n+7}_{\rm class\ AI}\left[\hat{P}\,({\bm{k}}_{\parallel})\right]=\nu^{d=8n+8}_{\rm class\ D}\left[\mathcal{H}_{\rm bulk}\,({\bm{k}})\right].$ (IV.15)
@@
Here, the zeroth Chern number ${\rm Ch}_{0}\,[\hat{P}$] is the rank of $\hat{P}$. These formulas (IV.5)-(IV.15) clarify the topology of the detached boundary states, leading to unique boundary responses and disorder effects.
@@
V Intrinsic non-Hermitian topology
[…]